# Automated QoR improvement in OpenROAD with coding agents

### v1

@@ Abstract.
EDA development and innovation has been constrained by scarcity of expert engineering resources. While leading LLMs have demonstrated excellent performance in coding and scientific reasoning tasks, their capacity to advance EDA technology itself has been largely untested. We present AuDoPEDA, an autonomous, repository-grounded coding system built atop OpenAI models and a Codex-class agent that reads OpenROAD, proposes research directions, expands them into implementation steps, and submits executable diffs. Our contributions include (i) a closed-loop LLM framework for EDA code changes; (ii) a task suite and evaluation protocol on OpenROAD for PPA-oriented improvements; and (iii) end-to-end demonstrations with minimal human oversight. Experiments in OpenROAD achieve routed wirelength reductions of up to 5.9%, and effective clock period reductions of up to 10.0%.
@@ 1. Introduction
Stage 3: Autonomous execution with verification loops. A Codex-class coding agent, operating in a planner–executor architecture, applies the proposed diffs, compiles the modified code, and runs instrumented OpenROAD flows on designated benchmarks. The agent iterates using self-measured signals (e.g., routed wirelength (rWL), worst negative slack (WNS), total negative slack (TNS), via count) together with guardrails such as revert-on-regression and bisect-on-failure. It leverages tool-use patterns (code search, syntax trees, build logs) similar to SWE-agent (sweagent, ), and can opportunistically invoke tool APIs learned from examples (cf. Toolformer) (toolformer, ). Failures (compile errors, runtime crashes, QoR regressions) are converted into counterexamples and fed back into the granular plan for repair, closing a self-correcting feedback loop.
@@
  • End-to-end validation on OpenROAD, demonstrating that the system can produce accepted patches and measurable QoR improvements — specifically, in routed wirelength and effective clock period — on fixed benchmarks and acceptance tests.
Scope and positioning. Unlike prior work that either documents code or assists local edits, our system couples (i) graph-structured documentation, (ii) literature-grounded planning, and (iii) agentic execution with QoR feedback, all tailored to an industrial-scale EDA repository. The design emphasizes programming LLMs rather than ad hoc prompt engineering: each stage is a typed operator with explicit inputs and outputs (graphs, cards, plans, diffs, metrics), enabling determinism, auditability, and reuse across flows (dspy, ). Below, Section 2 reviews relevant literature, and Section 3 presents the pipeline end-to-end. Section 4 gives full-system results: documentation artifacts, high-level and granular implementation plans, and resulting code diffs, along with QoR (routed wirelength, effective clock period) improvements obtained by applying these diffs to OpenROAD. Section 5 concludes with future directions for autonomous LLM agents in EDA.
@@ 3.4. S3: Autonomous Execution with QoR Feedback
  • measure: Extract PPA from tool reports (e.g., routed wirelength/DRCs from TritonRoute; WNS/TNS from OpenSTA).
@@ 5. Conclusions
We demonstrate that autonomous, repository-scale code modification for industrial-grade EDA flows is both feasible and effective when supported by structured documentation, literature-grounded planning, and agentic execution with QoR feedback. AuDoPEDA tightly couples graph-based documentation, DSPy-programmed research planning, localized granular plans, and a Codex-class executor to form an end-to-end system capable of generating legitimate, verifiable improvements to the OpenROAD physical design stack. Experiments across multiple PDKs, designs, and operating regimes show that the agent can (i) synthesize coherent research hypotheses, (ii) translate them into safe, diff-level implementations, and (iii) achieve measurable PPA improvements without per-design tuning or manual intervention. These include nontrivial ECP reductions under aggressive timing, and substantial routed wirelength improvements in high-utilization, congestion-heavy configurations. Importantly, all improvements are produced by a single autonomous diff per experiment category, underscoring AuDoPEDA’s ability to discover repository-scale, flow-consistent enhancements rather than one-off parameter tweaks.
Overall, AuDoPEDA suggests a new paradigm for EDA R&D via agents that are not mere assistants or script generators, but full participants that read code, propose algorithmic changes and validate their effects, and produce reproducible contributions. Future work includes exploring agentic code changes to improve design power metrics, integrating formal verification signals into the QoR gate, exploring multi-agent collaboration across flow stages, and extending the documentation and planning framework to additional open-source EDA systems. We believe that AuDoPEDA constitutes a first step toward self-improving, continuously learning design-automation toolchains.

### v2

@@ Abstract.
EDA development and innovation has been constrained by scarcity of expert engineering resources. While leading LLMs have demonstrated excellent performance in coding and scientific reasoning tasks, their capacity to advance EDA technology itself has been largely untested. We present AuDoPEDA, an autonomous, repository-grounded coding system built atop OpenAI models and a Codex-class agent that reads OpenROAD, proposes research directions, expands them into implementation steps, and submits executable diffs. Our contributions include (i) a closed-loop LLM framework for EDA code changes; (ii) a task suite and evaluation protocol on OpenROAD for PPA-oriented improvements; and (iii) end-to-end demonstrations with minimal human oversight. Experiments in OpenROAD achieve routed wirelength reductions of up to 5.9%, effective clock period reductions of up to 10.0%, and power reductions of up to 19.4%.
@@ 1. Introduction
Stage 3: Autonomous execution with verification loops. A Codex-class coding agent, operating in a planner–executor architecture, applies the proposed diffs, compiles the modified code, and runs instrumented OpenROAD flows on designated benchmarks. The agent iterates using self-measured signals (e.g., routed wirelength (rWL), worst negative slack (WNS), total negative slack (TNS), via count and power) together with guardrails such as revert-on-regression and bisect-on-failure. It leverages tool-use patterns (code search, syntax trees, build logs) similar to SWE-agent (sweagent, ), and can opportunistically invoke tool APIs learned from examples (cf. Toolformer) (toolformer, ). Failures (compile errors, runtime crashes, QoR regressions) are converted into counterexamples and fed back into the granular plan for repair, closing a self-correcting feedback loop.
@@
  • End-to-end validation on OpenROAD, demonstrating that the system can produce accepted patches and measurable QoR improvements — specifically, in routed wirelength, effective clock period, and total power consumption — across fixed benchmarks and acceptance tests.
Scope and positioning. Unlike prior work that either documents code or assists local edits, our system couples (i) graph-structured documentation, (ii) literature-grounded planning, and (iii) agentic execution with QoR feedback, all tailored to an industrial-scale EDA repository. The design emphasizes programming LLMs rather than ad hoc prompt engineering: each stage is a typed operator with explicit inputs and outputs (graphs, cards, plans, diffs, metrics), enabling determinism, auditability, and reuse across flows (dspy, ). Below, Section 2 reviews relevant literature, and Section 3 presents the pipeline end-to-end. Section 4 gives full-system results: documentation artifacts, high-level and granular implementation plans, and resulting code diffs, along with QoR (routed wirelength, effective clock period, and total power) improvements obtained by applying these diffs to OpenROAD. Section 5 concludes with future directions for autonomous LLM agents in EDA.
@@ 3.4. S3: Autonomous Execution with QoR Feedback
  • measure: Extract PPA from tool reports (e.g., routed wirelength/DRCs from TritonRoute; WNS, TNS, and total power from OpenSTA).
@@ 4.2. Better ECP at Aggressive Timing
4.3. Improvement of Power
For this experiment, we provide the system with the RSZ documentation and ask it to reduce power by modifying the resizer module in ORFS. The agent’s generated strategy autonomously develops a more advanced recovery flow than the baseline. Instead of only downsizing cells along slack-eligible timing paths, the strategy shifts to a design-wide instance view and applies staged recovery under timing safety checks. Its architecture, derived from the provided documentation and reflected in the modified RSZ code, is built on the following principal themes: First, it replaces path-by-path target selection with instance ranking based on power contribution and timing margin, then prioritizes high-power non-critical cells. Second, it extends the recovery flow beyond sizing by combining buffer removal, cell downsizing, and low-leakage VT swapping in a multi-pass sequence.
[…]
ASAP7 aes 153.74 150.75 1.947-1.947 9.836
ASAP7 ibex 58.10 47.24 18.703-18.703 2.606
ASAP7 jpeg 119.70 117.48 1.853-1.853 2.880
SKY130HD aes 456.94 437.55 4.245-4.245 3.529
SKY130HD ibex 93.20 75.14 19.383-19.383 4.433
SKY130HD jpeg 485.90 428.01 11.914-11.914 1.786
Nangate45 aes 385.29 373.67 3.017-3.017 1.339
Nangate45 ibex 95.98 89.37 6.883-6.883 3.644
Nangate45 jpeg 499.22 454.71 8.915-8.915 5.647
Table 5. Power outcomes (Pwr = total power, in mW) from power improvement testing.
Table 5 summarizes the experimental results, showing power reductions in all cases. [^4] All power values are extracted at the finish stage of the flow using OpenROAD report_power_metric, and the reported total power is the sum of internal, switching, and leakage power values. In addition, the RECOVER_POWER variable is set to 100 in the flow configuration of each design. OpenROAD/OpenSTA power and timing analyses have been previously reported to show a reasonable level of correlation with commercial tools (RDF_ICCAD24, ). This improvement stems from the difference in optimization scope. A path-based method concentrates its effort on specific timing paths, so it can miss high-power cells that reside on non-critical paths. In contrast, the instance-based method ranks all cells by their power contribution and available slack, then directly targets high-impact, non-critical instances first. As a result, it achieves larger power savings under the same timing constraints. In addition, buffer removal and VT swapping provide extra leverage beyond simple downsizing, reducing both dynamic power (via switching activity and capacitance) and leakage power, which further broadens the room for improvement. The last column reports the change in ECP, which captures the overhead induced by the new recovery flow. As expected, we observe ECP increases in all benchmarks (by roughly 1–10%), but this overhead is generally modest relative to the achieved power savings.
@@ 5. Conclusions
We demonstrate that autonomous, repository-scale code modification for industrial-grade EDA flows is both feasible and effective when supported by structured documentation, literature-grounded planning, and agentic execution with QoR feedback. AuDoPEDA tightly couples graph-based documentation, DSPy-programmed research planning, localized granular plans, and a Codex-class executor to form an end-to-end system capable of generating legitimate, verifiable improvements to the OpenROAD physical design stack. Experiments across multiple PDKs, designs, and operating regimes show that the agent can (i) synthesize coherent research hypotheses, (ii) translate them into safe, diff-level implementations, and (iii) achieve measurable PPA improvements without per-design tuning or manual intervention. These include nontrivial ECP reductions under aggressive timing, and substantial routed wirelength improvements in high-utilization, congestion-heavy configurations, and consistent power reductions across multiple PDKs and designs. Importantly, all improvements are produced by a single autonomous diff per experiment category, underscoring AuDoPEDA’s ability to discover repository-scale, flow-consistent enhancements rather than one-off parameter tweaks.
Overall, AuDoPEDA suggests a new paradigm for EDA R&D via agents that are not mere assistants or script generators, but full participants that read code, propose algorithmic changes and validate their effects, and produce reproducible contributions. Future work includes integrating formal verification signals into the QoR gate, exploring multi-agent collaboration across flow stages, and extending the documentation and planning framework to additional open-source EDA systems. We believe that AuDoPEDA constitutes a first step toward self-improving, continuously learning design-automation toolchains.
Acknowledgments
This work is partially supported by the Samsung AI Center.